# Particle motion around luminous neutron stars: effects of deviation from Schwarzschild spacetime

### v1

@@ Abstract
We study trajectories of test particles around a luminous, static, spherically symmetric neutron star, under the combined influence of gravity and radiation. In general relativity, for Schwarzschild spacetime, an equilibrium sphere (the Eddington Capture Sphere) is formed for near-Eddington luminosities. We generalize these results to a broad class of static, spherical spacetimes. We also study the dynamics of particles in a strong radiation field in spherical spacetimes. The results are illustrated for two cases, Reissner-Nordström spacetime of a charged spherical object in general relativity and Kehagias-Sfetsos spacetime, arising from the Hořava-Lifshitz gravity theory. Our findings apply to neutron stars under gravitational field equations different from the vacuum Einstein field equations of general relativity, such as in modified theories of gravity, the only requirement being that test particles follow geodesics in the absence of the radiation field. We show that it is possible, in principle, to test the gravity theory by investigating such test particle equilibria and trajectories with observations of neutron stars’ X-ray bursts.
@@ I Introduction
results in a global equilibrium between the Newtonian gravity and the radiation pressure for a spherical source of mass $M$. In Eq. (1) we assume hydrogen plasma, hence $\sigma_{\rm T}$ is the Thomson cross section and $m$ is the proton mass. $G$ and $c$ denote the gravitational constant and speed of light, respectively. When it is not confusing, we use $G=c=1$ units hereafter. The problem of the gravity-radiation equilibrium is more complicated for the relativistic case – in general relativity (GR) equilibrium can only be established at a particular radius, referred to as the Eddington Capture Sphere (ECS; see Stahl2012 ; Wielgus2012 ). The ECS radius $r_{\rm ECS}$ depends on the stellar luminosity as seen by a distant observer $L_{\infty}$, and in Schwarzschild spacetime corresponds to
@@
where $\lambda\equiv L_{\infty}/L_{\rm Edd}$. In Eq. (2), $L_{\infty}=L_{\rm{Edd}}$ represents equilibrium established at infinity. The ECS may form as a transient phenomenon during type-I luminous X-ray bursts on neutron stars Tawara1984 ; Lewin1993 , particularly in cases exhibiting photospheric expansion Lewin1984 ; Strohmayer2006 . Studying such events may lead to constraints on neutron star equation of state Ozel2006 ; Lattimer2012 , mass, radius Galloway2008 ; Bollimpalli2019 , and spin Muno2001 . In this paper we consider whether the location of the equilibrium surface and test-particle dynamics are sensitive to deviations of the underlying metric from the Schwarzschild solution. While several authors considered deviations caused by the central mass rotation SokOh2010 ; Bini2011 ; DeFalco2019 ; Wielgus2019 , quadrupolar deviation of the metric from spherical symmetry DeFalco2020 , or both DeFalco2021 , we focus on the effects related to modifying the underlying gravity field equations (and/or modifications of the background metric).
Furthermore, the aim of this paper is to demonstrate the impact of the spacetime deviation from the canonical Schwarzschild case on properties of systems including luminous neutron stars, illustrating it with concrete examples. Such deviations in spacetime geometry could be in principle detected through detailed modeling of X-ray light curves of bursting neutron stars, offering a previously unexplored avenue to test theories of gravity in the strong field and large curvature regime.
@@ II Equations of motion
where $a^{\mu}$ is the particle’s four-acceleration and $u^{\mu}$ is its four-velocity. The only non-zero Christoffel symbols are
@@
Using the projection tensor $h^{\mu}_{\ \nu}$ onto the particle’s local 3-space, Kronecker tensor $\delta^{\mu}_{\ \nu}$, and the radiation stress-energy tensor $T^{\nu\rho}$, the radiation flux $F^{\mu}$ measured in the particle’s local rest frame can be represented as
@@
Figure 1: Radius of gravity-radiation equilibrium sphere for Reissner-Nordström (top row) and Kehagias-Sfetsos (bottom row) spacetimes for various luminosity parameters $\lambda$. Left column: radius as function of the metric parameter $(Q/M)^{2}$ in RN spacetime ($\omega M^{2}$ in KS spacetime), $Q=0$ ($\omega M^{2}\rightarrow\infty$) corresponds to Schwarzschild spacetime. Right column: relative change in the location of the equilibrium sphere with respect to the Schwarzschild case.
Figure 2: Radius (left column), radial velocity component $v^{r}=\sqrt{-g_{rr}/g_{tt}}\,u^{r}/u^{t}$ (middle column), and energy $u_{t}$ (right column) as functions of the proper time $\tau$ for particles ejected from the neutron star surface located at $R=4M$ by the instantaneous burst of radiation with the luminosity parameter $\lambda=0.75$ (solid), $\lambda=0.80$ (dashed), and $\lambda=0.85$ (dotted). The black lines correspond to Schwarzschild spacetime. The red lines in the top row correspond to Reissner-Nordström spacetime with $(Q/M)^{2}=1$. The red lines in the bottom row correspond to Kehagias-Sfetsos spacetime with $\omega M^{2}=0.5$.
@@
where $M$ in a general spacetime corresponds to the ADM mass, so we assume that the spacetime approaches Schwarzschild solution in the weak field (large radius) limit. Putting together Eqs. (4)–(20) and leveraging spherical symmetry to choose a coordinate system with $\theta=\pi/2$, ${\rm d}\theta/{\rm d}\tau={\rm d}^{2}\theta/{\rm d}\tau^{2}=0$, we obtain the full equations of motion for a test particle
@@
There are four components in Eq. (II), that can be associated with gravitational acceleration $\propto\xi^{\prime}$, centrifugal acceleration $\propto({\rm d}\phi/{\rm d}\tau)^{2}$, radiation pressure $\propto\lambda\,u^{t}$, and radial radiation drag $\propto\lambda\,{\rm d}r/{\rm d}\tau$. In Eq. (22) we recognize the Coriolis acceleration $\propto({\rm d}r/{\rm d}\tau)({\rm d}\phi/{\rm d}\tau)$. The second component of Eq. (22) is the Poynting–Robertson drag force, removing particle’s angular momentum through interaction with the radiation field. In general relativity these effects were studied in, e.g., Bini2009 ; Stahl2013 ; Mishra2014 ; DeFalco2019 ; Fragile2020 . Equations (II)–(23) reduce to the ones presented in Stahl2012 for the case of Schwarzschild spacetime.
II.1 Radial equilibrium surface
@@
If Eq. (26) has a solution for some $r=r_{0}>R$, then it physically corresponds to the gravity-radiation equilibrium surface suspended above the luminous stellar surface. For Schwarzschild metric Eq. (26) reduces to $\lambda=\xi^{1/2}$, and hence the ECS formula given in Eq. (2) is recovered. We denote this general equilibrium radius with $r_{0}$, and the Schwarzschild value $r_{\rm ECS}\equiv r_{0,\rm{Schw}}$.
@@ III Metric examples
We exemplify the general calculations given in Section II by considering two particular spacetime models: Reissner-Nordström (RN) and Kehagias-Sfetsos (KS; kehagiasSfetsos2009PhLB, ). The former originates from the electrovacuum Einstein field equations and corresponds to the spacetime of a static electrically charged black hole (or exterior of a charged spherical star). Recently a lot of interest in RN spacetime came from the brane cosmology Randall1999 , where the RN model is a solution of the effective field equation and the “tidal charge” corresponds to gravitational effects from the fifth dimension Dadhich2000 . The KS spacetime arises as a static, spherical solution representing a black hole (or exterior of a star) in Horava’s gravity, a popular model of a quantum field theory of gravity Horava2009 . The geodesic structure of the KS spacetime has been studied extensively by katkaVieiraEtal2015GRG ; stuchlikSchee2014CQG ; vieiraMarekEtal2014PRD , for instance.
@@ III.1 Equilibrium spheres
Figure 4: Hoyle-Lyttleton accretion onto a luminous star. The particles are initiated at horizontal axis coordinate $x=-5000M$, where the gravitational and radiative impact of the star is negligible, all with initial velocity of $v=0.1\,c$ along the $x$ direction. The star is characterized by $R=4M$ and $\lambda=0.99$. Left: Schwarzschild spacetime. Trajectories with relatively low impact parameter are captured by the equilibrium sphere, whereas trajectories with high impact parameter are scattered (to infinity). The red curve corresponds to the lowest impact parameter for scattered trajectories (to infinity) considering a $1M$ resolution in the grid of impact parameters, $b_{\rm crit}=1697M$. Middle: RN spacetime with $(Q/M)^{2}=1$. The red curve curve corresponds to the lowest impact parameter for scattered trajectories $b_{\rm crit}=1615M$. Typical captured trajectories in the middle panel penetrate the equilibrium sphere and then bounce, crossing it outwards and finally settling into it. The same qualitative picture happens in the right panel (KS spacetime with $\omega M^{2}=0.5$), but with a much smaller bouncing amplitude and with $b_{\rm crit}=1697M$, resembling the behavior in the Schwarzschild spacetime.
@@ III.2 Particle ejection
We study the scenario of a luminous outburst, expelling particles from the neutron star surface. Under the assumption of spherical symmetry, this becomes a one-dimensional problem for the particle’s radial motion. Similar problems were considered for the Schwarzschild case, e.g., by Wielgus2012 ; Stahl2013 ; Mishra2014 , including more general configurations. Here, we consider only the most physically relevant case, corresponding to particles ejected (with zero initial velocity) from the neutron star surface by a sudden burst of luminosity to a near-Eddington value. This is essentially a toy model for the photospheric radius expansion X-ray bursts of neutron stars Lewin1984 ; Lewin1993 ; Kuulkers2003 . Example results are shown in Fig. 2. The timescale for the particles to settle on the equilibrium sphere is no longer than 1000 $M$ for the chosen peak luminosity $\lambda$, which is of order of 0.01 s for a typical neutron star. The radial velocity of ejected particles may be significantly larger for non-Schwarzschild spacetimes, reaching around 0.1 $c$ for RN spacetime, before the particle settles on the equilibrium sphere. The kinetic energy of the particle is removed by the radiative drag force (Fig. 2, second column), but the particle energy $u_{t}$ is not conserved and grows as a result of the radiation pressure (Fig. 2, last column).
@@ III.3 Hoyle-Lyttleton accretion
In Fig. 4 we present trajectories of particles in the Hoyle-Lyttleton scenario for Schwarzschild (left), extreme RN (middle), and extreme KS (right) spacetimes. The setup corresponds to a collection of particles approaching the neutron star from the left with an initial velocity of 0.1 $c$, set beyond the gravitational sphere of influence of the neutron star. The particles are parameterized by their impact parameter, proportional to their (initial) specific angular momentum with respect to the neutron star. Particles with impact parameter $b$ lower than certain critical value $b_{\rm crit}$ (specific angular momentum $\ell<\ell_{\rm crit}$) are captured by the equilibrium sphere, hence the name Eddington Capture Sphere. The value of the $b_{\rm crit}$ parameter (and therefore of $\ell_{\rm crit}$) depends on the initial velocity, and also (weakly) on the neutron star diameter, impacting dynamical radiation drag forces and hence the efficiency of the angular momentum removal. These effects were investigated by, among others, Bini2009 ; Stahl2012 for the Schwarzschild spacetime case and by SokOh2011 ; Wielgus2019 for a slowly rotating neutron star. Here we demonstrate that the role of the spacetime geometry may also be significant, which is particularly clear for the (extreme) RN case in Fig. 4. It is interesting to observe that (extreme) RN spacetime, with the equilibrium sphere larger than those of Schwarzschild and (extreme) KS cases, actually corresponds to a smaller cross section for capturing particles. This highlights the dynamic character of this problem, and the role of the relativistic radiation drag forces acting on a moving particle and depending on its four-velocity.

### v2

@@ Abstract
We study trajectories of test particles around a luminous, static, spherically symmetric neutron star, under the combined influence of gravity and radiation. In general relativity, for Schwarzschild spacetime, an equilibrium sphere (the Eddington Capture Sphere) is formed for near-Eddington luminosities. We generalize these results to a broad class of static, spherical spacetimes. We also study the dynamics of particles in a strong radiation field in spherical spacetimes. The results are illustrated for two cases, Reissner-Nordström spacetime of a charged spherical object in general relativity and Kehagias-Sfetsos spacetime, arising from the Hořava–Lifshitz gravity theory. Our findings apply to neutron stars under gravitational field equations different from the vacuum Einstein field equations of general relativity, such as in modified theories of gravity, the only requirement being that test particles follow geodesics in the absence of the radiation field. The effects that we describe are, in principle, measurable through observations of X-ray bursts of neutron stars. Hence, detailed future studies could use such observations to test gravity theories in the strong-field regime, provided that the impact of the spacetime geometry can be disentangled from the astrophysical uncertainties.
@@ I Introduction
results in a global equilibrium between the Newtonian gravity and the radiation pressure for a spherical source of mass $M$. In Eq. (1) we assume hydrogen plasma, hence $\sigma_{\rm T}$ is the Thomson cross section and $m$ is the proton mass. $G$ and $c$ denote the gravitational constant and speed of light, respectively. The problem of the gravity-radiation equilibrium is more complicated for the relativistic case – in general relativity (GR) equilibrium can only be established at a particular radius, referred to as the Eddington Capture Sphere (ECS; see Stahl2012 ; Wielgus2012 ). The ECS radius $r_{\rm ECS}$ depends on the stellar luminosity as seen by a distant observer $L_{\infty}$, and in Schwarzschild spacetime corresponds to
@@
where $\lambda\equiv L_{\infty}/L_{\rm Edd}$. In Eq. (2), $L_{\infty}=L_{\rm{Edd}}$ represents equilibrium established at infinity. The ECS may form as a transient phenomenon during type-I luminous X-ray bursts on neutron stars Tawara1984 ; Lewin1993 , particularly in cases exhibiting photospheric expansion Lewin1984 ; Strohmayer2006 , when the radiation originating from a thermonuclear explosion engulfing the neutron star surface temporarily balances the effective gravity. Studying such events may lead to constraints on neutron star equation of state Ozel2006 ; Lattimer2012 , mass, radius Galloway2008 ; Bollimpalli2019 , and spin Muno2001 . In this paper we consider whether the location of the equilibrium surface and test-particle dynamics are sensitive to deviations of the underlying metric from the Schwarzschild solution. While several authors considered deviations caused by the central mass rotation SokOh2010 ; Bini2011 ; DeFalco2019 ; Wielgus2019 , quadrupolar deviation of the metric from spherical symmetry DeFalco2020 , or both DeFalco2021 , we focus on the effects related to modifying the underlying gravity field equations (and/or modifications of the background metric).
Furthermore, the aim of this paper is to demonstrate the impact of the spacetime deviation from the canonical Schwarzschild case on properties of systems including luminous neutron stars, illustrating it with concrete examples. Such deviations in spacetime geometry could in principle be detected through detailed modeling of X-ray light curves of bursting neutron stars, offering a previously unexplored avenue to test theories of gravity in the strong field and large curvature regime. This task will, however, require careful modeling in order to disentangle the effects of astrophysical configuration, star rotation, and potential deviation from general relativity.
@@ II Equations of motion
where $a^{\mu}$ is the particle’s four-acceleration and $u^{\mu}$ is its four-velocity. Equation (4) corresponds to the covariant formulation of Newton’s second law with a standard electrodynamic radiation force term $f^{\mu}\propto F^{\mu}$, proportional to the radiation flux measured in the frame of the test particle. The only non-zero Christoffel symbols are
@@
Using the projection tensor $h^{\mu}_{\ \nu}$ onto the particle’s local 3-space, Kronecker tensor $\delta^{\mu}_{\ \nu}$, and the radiation stress-energy tensor $T^{\nu\rho}$, the radiation flux $F^{\mu}$ measured in the particle’s local rest frame can be represented as  [^1] Unless explicitly stated, we hereafter use geometrized units ($G=c=1$).
@@
Figure 1: Radius of the gravity-radiation equilibrium sphere for Reissner-Nordström (top row) and Kehagias-Sfetsos (bottom row) spacetimes for various luminosity parameters $\lambda$. Left column: radius as function of the metric parameter $(Q/M)^{2}$ in RN spacetime ($\omega M^{2}$ in KS spacetime), $Q=0$ ($\omega M^{2}\rightarrow\infty$) corresponds to Schwarzschild spacetime. Right column: relative change in the location of the equilibrium sphere with respect to the Schwarzschild case.
Figure 2: Radius (left column), radial velocity component $v^{r}=\sqrt{-g_{rr}/g_{tt}}\,u^{r}/u^{t}$ (middle column), and energy $-u_{t}$ (right column) as functions of the proper time $\tau$ for particles ejected from the neutron star surface located at $R=4M$ by the instantaneous burst of radiation with the luminosity parameter $\lambda=0.75$ (solid), $\lambda=0.80$ (dashed), and $\lambda=0.85$ (dotted). The time unit corresponds to $M\approx 10^{-5}$ s for a neutron star mass of $2\,M_{\odot}$. The black curves correspond to Schwarzschild spacetime. The red curves in the top row correspond to Reissner-Nordström spacetime with $(Q/M)^{2}=1$. The red curves in the bottom row correspond to Kehagias-Sfetsos spacetime with $\omega M^{2}=0.5$.
@@
where $M$ in a general spacetime corresponds to the Arnowitt-Deser-Misner mass, so we assume that the spacetime approaches Schwarzschild solution in the weak field limit (large radius). Putting together Eqs. (4)–(20) and leveraging spherical symmetry to choose a coordinate system with $\theta=\pi/2$, ${\rm d}\theta/{\rm d}\tau={\rm d}^{2}\theta/{\rm d}\tau^{2}=0$, we obtain the full equations of motion for a test particle
@@
There are four components in Eq. (II), that can be associated with gravitational acceleration $\propto\xi^{\prime}$, centrifugal acceleration $\propto({\rm d}\phi/{\rm d}\tau)^{2}$, radiation pressure $\propto\lambda\,u^{t}$, and radial radiation drag $\propto\lambda\,{\rm d}r/{\rm d}\tau$. In Eq. (22) we recognize the Coriolis acceleration $\propto({\rm d}r/{\rm d}\tau)({\rm d}\phi/{\rm d}\tau)$. The second component of Eq. (22) is the Poynting–Robertson drag force, removing the particle’s angular momentum through interaction with the radiation field. In general relativity these effects were studied in, e.g., Bini2009 ; Stahl2013 ; Mishra2014 ; DeFalco2019 ; Fragile2020 . Equations (II)–(23) reduce to the ones presented in Stahl2012 for the case of Schwarzschild spacetime.
II.1 Back-reaction on spacetime
Another interesting question that one may ask is whether the strong radiation field, balancing the effective gravity, should be included in the calculation of the spacetime metric from the underlying field equations as a non-zero stress-energy tensor component. While a complete, rigorous and non-linear treatment of this problem is complicated even for fixed underlying field equations, one can make a rather simple order-of-magnitude argument following Section 4 of (Wielgus2014, ), where a very similar problem of radiation field back-reaction on the Schwarzschild spacetime geometry was considered. Assuming Einstein vacuum field equations we characterize the magnitude of the Ricci part of the Riemann tensor by computing a scalar quantity
$\mathcal{R}=R^{\mu\nu}R_{\mu\nu}=\kappa^{2}T^{\mu\nu}T_{\mu\nu}\propto\frac{(\lambda L_{\rm Edd})^{2}}{r^{4}}$ (25)
with $\kappa=8\pi G/c^{4}$. After some amount of algebra and taking liberal overestimates of various quantities such as stellar radius $R$ or $\xi_{R}$ we conclude that
$\mathcal{R}<10^{-40}\frac{c^{8}}{G^{4}M^{4}}$ (26)
near the surface of the Eddington-luminosity neutron star. This value characterizes the impact of the stress-energy tensor $T^{\mu\nu}$ on the spacetime geometry and can be compared with the Weyl part $C^{abcd}$ of the Riemann tensor $R^{abcd}$, characterizing the magnitude of the spacetime curvature in the absence of $T^{\mu\nu}$. The latter can be evaluated for the vacuum background metric simply with the Kretschmann scalar, which characterizes the Weyl tensor magnitude for a vacuum background spacetime
$\mathcal{K}$ $=R^{abcd}R_{abcd}=C^{abcd}C_{abcd}$
$=\frac{48}{r^{6}}\frac{G^{2}M^{2}}{c^{4}}>10^{-11}\frac{c^{8}}{G^{4}M^{4}}$ (27)
where we only assumed that $r<100\,GM/c^{2}$. While this example is limited to Schwarzschild background and only constitutes a linear approximation, as the stress-energy tensor is evaluated in the background spacetime without accounting for the back-reaction, the huge gap between the magnitudes of $\mathcal{K}$ and $\mathcal{R}$ shows that the back-reaction is negligible for the Eddington-luminosity radiation around a neutron star. As a matter of fact, it would take over 10 orders of magnitude increase of the stellar luminosity for the energy of radiation to start contributing to the spacetime curvature in an appreciable way, as we show that $\mathcal{R}/\mathcal{K}<10^{-29}$ and $\mathcal{R}\propto\lambda^{2}$.
II.2 Radial equilibrium surface
@@
If Eq. (29) has a solution for some $r=r_{0}>R$, then it physically corresponds to the gravity-radiation equilibrium surface suspended above the luminous stellar surface. For Schwarzschild metric Eq. (29) reduces to $\lambda=\xi^{1/2}$, and hence the ECS formula given in Eq. (2) is recovered. We denote this general equilibrium radius by $r_{0}$, and the Schwarzschild value $r_{\rm ECS}\equiv r_{0,\rm{Schw}}$.
@@ III Metric examples
We exemplify the general calculations given in Section II by considering two particular spacetime models: Reissner-Nordström (RN) and Kehagias-Sfetsos (KS; kehagiasSfetsos2009PhLB, ). The former originates from the electrovacuum Einstein field equations and corresponds to the spacetime of a static electrically charged black hole (or exterior of a charged spherical star). Recently a lot of interest in RN spacetime came from the brane cosmology Randall1999 , where the RN model is a solution of the effective field equations and the “tidal charge” corresponds to gravitational effects from the fifth dimension Dadhich2000 . The KS spacetime arises as a static, spherical solution representing a black hole (or the exterior of a star) in Horava’s gravity, a popular model of a quantum field theory of gravity Horava2009 . The geodesic structure of the KS spacetime has been studied extensively by katkaVieiraEtal2015GRG ; stuchlikSchee2014CQG ; vieiraMarekEtal2014PRD , for instance.
@@ III.1 Equilibrium spheres
Figure 4: Hoyle-Lyttleton accretion onto a luminous star. The particles are initiated at horizontal axis coordinate $x=-5000M$, where the gravitational and radiative impact of the star is negligible, all with initial velocity of $v=0.1\,c$ along the $x$ direction. The star is characterized by $R=4M$ and $\lambda=0.99$. Left: Schwarzschild spacetime. Trajectories with relatively low impact parameter are captured by the equilibrium sphere, whereas trajectories with high impact parameter are scattered (to infinity). The red curve corresponds to the lowest impact parameter for scattered trajectories (to infinity) considering a $1M$ resolution in the grid of impact parameters, $b_{\rm crit}=1697M$. Middle: RN spacetime with $(Q/M)^{2}=1$. The red curve curve corresponds to the lowest impact parameter for scattered trajectories $b_{\rm crit}=1615M$. Typical captured trajectories in the middle panel penetrate the equilibrium sphere and then bounce, crossing it outwards and finally settling into it. The same qualitative picture happens in the right panel (KS spacetime with $\omega M^{2}=0.5$), but with a much smaller bouncing amplitude and with $b_{\rm crit}=1697M$, resembling the behavior in Schwarzschild spacetime.
@@ III.2 Particle ejection
We study the scenario of a luminous outburst, expelling particles from the neutron star surface. Under the assumption of spherical symmetry, this becomes a one-dimensional problem for the particle’s radial motion. Similar problems were considered for the Schwarzschild case, e.g., by Wielgus2012 ; Stahl2013 ; Mishra2014 , including more general configurations. Here, we consider only the most physically relevant case, corresponding to particles ejected (with zero initial velocity) from the neutron star surface by a sudden burst of luminosity to a near-Eddington value. This is essentially a toy model for the photospheric radius expansion X-ray bursts of neutron stars Lewin1984 ; Lewin1993 ; Kuulkers2003 . Example results are shown in Fig. 2. The timescale for the particles to settle on the equilibrium sphere is no longer than 1000 $M$ for the chosen peak luminosity $\lambda$, which is of order of 0.01 s for a typical neutron star. The radial velocity of ejected particles may be significantly larger for non-Schwarzschild spacetimes, reaching around 0.1 $c$ for RN spacetime, before the particle settles on the equilibrium sphere. The kinetic energy of the particle is removed by the radiative drag force (Fig. 2, second column), but the particle energy $(-u_{t})$ is not conserved and grows as a result of the radiation pressure (Fig. 2, last column).
@@ III.3 Hoyle-Lyttleton accretion
In Fig. 4 we present trajectories of particles in the Hoyle-Lyttleton scenario for Schwarzschild (left), extreme RN (middle), and extreme KS (right) spacetimes. The setup corresponds to a collection of particles approaching the neutron star from the left with an initial velocity of 0.1 $c$, set beyond the gravitational sphere of influence of the neutron star. The particles are parameterized by their impact parameter, proportional to their (initial) specific angular momentum with respect to the neutron star. Particles with impact parameter $b$ lower than certain critical value $b_{\rm crit}$ (specific angular momentum $\ell<\ell_{\rm crit}$) are captured by the equilibrium sphere, hence the name Eddington Capture Sphere. The value of the $b_{\rm crit}$ parameter (and therefore of $\ell_{\rm crit}$) depends on the initial velocity, and also (weakly) on the neutron star diameter, impacting dynamical radiation drag forces and hence the efficiency of the angular momentum removal. Thus, the rate of mass capture, proportional to the capture cross section $\pi b_{\rm crit}^{2}$ of the luminous star, particle velocity, and particle mass density, is also sensitive to the spacetime geometry. These effects were investigated by, among others, Bini2009 ; Stahl2012 ; Wielgus2012 for the Schwarzschild spacetime case and by SokOh2011 ; Wielgus2019 for a slowly rotating neutron star. Here we demonstrate that the role of the spacetime geometry may also be significant, which is particularly clear for the (extreme) RN case in Fig. 4. It is interesting to observe that (extreme) RN spacetime, with the equilibrium sphere larger than those of Schwarzschild and (extreme) KS cases, actually corresponds to a smaller cross section for capturing particles. This highlights the dynamic character of this problem, and the role of the relativistic radiation drag forces acting on a moving particle and depending on its four-velocity.
Luminous type-I neutron star bursts are transient phenomena of thermonuclear origin, fueled by the ignited material accreting onto the stellar surface. The bursts occur on timescales much longer than the dynamical ones (since $GM_{\odot}/c^{3}\approx 5\times 10^{-6}$ s, the dynamical timescales involved are of order of milliseconds) and near-Eddington luminosity can be sustained for a duration $\Delta T$ of tens of seconds (Lewin1993, ), while the ignited fuel reservoir burns away. Hence, the total mass of the material accumulated on the equilibrium sphere always remains negligible in comparison to the stellar mass $M$ for physically motivated parameters. For the example parameters considered in Fig. 4 we find ($v=0.1\,c$, $b_{\rm crit}\approx 1700\,M$, $m_{p}$ is the proton mass, $n_{p}$ is the number density of particles)
$\frac{M_{\rm shell}}{M}$ $=\pi b^{2}_{\rm crit}m_{p}n_{p}v\frac{\Delta T}{M}$ (36)
$\approx 9\times 10^{-17}\left(\frac{\Delta T}{100\,\rm{s}}\right)\left(\frac{2M_{\odot}}{M}\right)\left(\frac{n_{p}}{10^{6}{\rm cm}^{-3}}\right)\,.$
The reference number density of $10^{6}\,{\rm cm}^{-3}$ overestimates a typical number density of molecular clouds.